# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ II.1 Reduced stochastic models: theoretical and practical considerations
  • Coarse-grained dynamics. In spatiotemporal systems like climate, coarse-graining involves: (i) selecting a few relevant modes, (ii) isolating specific temporal scales, and (iii) choosing a limited set of variables. Regarding point (i), we reduce the dimensionality of the system in terms of some bases, a common choice is the one of Empirical Orthogonal Functions (EOFs) [39, 8]. We retain only the first $n<N$ of such bases, and project the data onto them to obtain the principal components (PCs), which serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as those in [64, 62, 65] (among many others) may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [66, 67]. Considering the “right” coarse-grained procedures is perhaps the most difficult step in modeling from data. There is in fact no general mathematical guideline in high-dimensional dynamical systems for choosing the proper variables to model, an old problem [68] emphasized in many new contributions, such as [31, 69, 47, 48].
  • Inferring a stochastic model. Given a reduced order representation of the system, we are now left with the task of inferring a stochastic model of the form of Eq. (1). Many studies have shown that, for specific observables (e.g., sea surface temperature anomalies), a linear drift with additive noise provides forecasting skill comparable to full complexity models [4, 5, 70, 71]. This leads to a Linear Inverse Model (LIM) [3] of the form
    $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W},$
    where $\mathbf{L}$ is a linear operator, $\mathbf{\Sigma}$ is a noise matrix and $\mathbf{W}$ is the standard Wiener process. An interesting recent extension to tackle colored noise can be found in [72]. The multilevel regression proposed by Kratsov et al. (2005) [6, 7] generalizes these ideas by incorporating quadratic nonlinearities and non-Markovian effects with a series of “matryoshka”-like levels for residuals. In many practical cases, however, temporal coarse-graining (e.g., averaging over a few time steps) can effectively remove memory effects, allowing a Markovian closure [73]. In the framework proposed in [6, 7], the drift consists of both linear and quadratic nonlinear components. The inclusion of quadratic nonlinearities is physically motivated rather than ad-hoc. As shown in Section 2 of Lorenz (1963) [74], projecting forced-dissipative flows (in periodic domains) onto Fourier modes, yield a set of infinite coupled ODEs with linear and quadratic interactions for the modes’ amplitudes. In particular, the quadratic nonlinearity $\mathbf{B}(\mathbf{x},\mathbf{x})$ should be energy-conserving, satisfying $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$. This abstract representation underlies the starting point of the Majda-Timofeyev-Vanden Eijnden (MTV) mode reduction framework [49, 50, 57, 58]. Additionally, Majda and Harlim focused on the energy conservation constraint above to propose physically constrained regression models [11]. Guided by these insights, the neural models proposed in this work will have a deterministic dynamics specified by a linear term $\mathbf{L}$ and general nonlinear terms $\mathbf{n}(\mathbf{x})$; in formulae:
    $\frac{d\mathbf{x}}{dt}=\mathbf{L}\mathbf{x}+\mathbf{n}(\mathbf{x}).$
    It is in fact a sensible choice to guide the learning by specifying a decomposition into linear and nonlinear terms, instead of letting the neural network “figure out” the functional form by itself (see also [75]). Furthermore, we will move beyond additive noise and incorporate multiplicative (state-dependent) noise, both in one and higher dimensions. As we will show, incorporating state-dependent noise can be at times critical for accurately capturing the system’s response to external perturbations in the presence of partial observations.
@@ II.2 Response theory: a (very) brief overview
In this paper, we argue that a reduced-order model should capture not only the system’s stationary statistics but also its response to external perturbations. This naturally raises the question: How to assess skillful responses to general perturbations? Since it is impossible to test responses to the full space of possible forcings, linear response theory provides a principled framework for analyzing the system’s linear response to perturbations.
@@
[…]
@@
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77]. The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$; under the assumption that the reduced order models correctly capture the physics of the system. For more details to this physics-based approach to causal inference see [78, 79, 47, 80, 62].
@@
A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom. Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation, see e.g. discussion in Falasca et al. (2024) [62]. Since the linear response to small perturbations is given by the convolution formula in Eq. (4), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings. For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to reliably capture at least the linear component of responses to general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well.
Inferring the response operator in practice.
  1. 1.
    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient. We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor. $N_{e}$ should be very large (i.e. $N_{e}\gg 1$) in order to sample the whole attractor.
  2. 2.
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $x_{j}(0)\rightarrow x_{j}(0)+\Delta_{j}\delta(t)$ of size $\Delta_{j}$ on the $j$-th degree of freedom at time $t=0$. $\delta(t)$ is the Dirac delta function. The amplitude $\Delta_{j}$ should be theoretically infinitesimally small. In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
  3. 3.
  4. 4.
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$. The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the perturbed distribution, therefore quantifying the response in ensemble mean and variance.
  5. 5.
    We define the impulse response operator for observable $A(x_{k}(t))$ as
    $R_{k,j}(t)=\frac{\delta A(x_{k}(t))}{\Delta_{j}}=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
As argued by Cecconi et al. (2012) [31], the main challenge in modeling real-world dynamical systems from data stems from their high dimensionality. Learning a dynamical system from data requires sufficiently long time series to sample many recurrences. However, the recurrence time, i.e. the time needed for a trajectory to return to a previous region in state space, scales exponentially with the system’s dimensionality [31, 81]. The success of data-driven methods in real-world systems is therefore often attributed to the presence of an effective low-dimensional dynamics. In other words, the dynamics of the system usually lives on a low-dimensional manifold embedded in a higher-dimensional space. This so-called “manifold hypothesis” underpins research in fields ranging from spatiotemporal chaos [82, 83, 84, 85] to neuroscience [86, 87, 88, 89, 90], biophysics [91, 92], artificial neural networks [93, 94] and climate dynamics [74, 1, 95, 96, 54]. In practice, however, even when a system exhibits low effective dimensionality, we still face several difficult challenges. First, it is often unclear how to construct an optimal basis for efficient modeling in the system’s intrinsic coordinates [64]. Second, and more fundamentally, we typically observe only a subset of relevant variables of a complex, multiscale physical system. Here we illustrate how learning on only partial observations, as is common in real-world applications [10, 33], can hinder data-driven modeling, even when the underlying dynamics is low-dimensional and seemingly tractable. Importantly, while the discussion is framed in terms of neural emulators, motivated by current challenges in climate science, the issues we highlight are general and inherent to data-driven approaches for real-world dynamical systems.
III.1 A triad model as a prototype stochastic climate model
We consider a triad model originally introduced by Majda and collaborators in the context of low-frequency climate dynamics [34, 36]. The model is:
$\begin{split}dx_{1}&=(L_{11}x_{1}+L_{12}x_{2}+L_{13}x_{3}+Ix_{1}x_{2})dt\\ dx_{2}&=(-L_{12}x_{1}-Ix_{1}^{2}-\gamma_{2}\epsilon^{-1}x_{2})dt+\sigma_{2}\epsilon^{-1/2}dW_{2}\\ dx_{3}&=(-L_{13}x_{1}-\gamma_{3}\epsilon^{-1}x_{3})dt+\sigma_{3}\epsilon^{-1/2}dW_{3}\\ \end{split}$ (5)
@@ III.1 A triad model as a prototype stochastic climate model
This idealized system mimics the functional structure of a reduced stochastic climate model, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities. The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics. The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting. The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations. This stochastic forcing, together with the linear damping terms $\gamma_{2}$ and $\gamma_{3}$, serves as a linear proxy for the nonlinear self-interactions of the fast variables (see Eq. (4) in [49] or Section 3.3 in [37] for further details). Importantly, the model generates relatively small, but significant, deviations from Gaussianity which is a relevant feature in the study of low-frequency climate dynamics [34] and found in both general circulation models and observational data, see e.g. [97, 98, 99].
III.2 Data-driven modeling when we observe the full state vector
@@
We consider the relatively simple case of a low-dimensional dynamical system with only one characteristic time scale (i.e. $\epsilon=1$). The parameters chosen for this first test are: $L_{11}=-2$, $L_{12}=1$, $L_{13}=1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=1$. We now integrate the system in (5) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units). As we will show later, given the system’s decorrelation time scale, the resulting time series can be effectively treated as infinitely long. With this long trajectory of the full state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$ in hand, we now attempt to learn data-driven models directly from the time series. We construct both linear and nonlinear stochastic models. We specify the functional forms of these models and assess their ability to reproduce both the stationary and perturbed statistics of the original system defined in Equation (5).
@@ III.2.1 Functional form of the data-driven models considered
We consider two data-driven models and report their functional forms below. To keep the main text focused, we leave the technical details of the fitting procedures, along with details of the neural networks, in Appendix A.
@@
    $\mathbf{L}$ is a linear operator and $\mathbf{n}(\mathbf{x})$ is parametrized by a Multilayer Perceptron (MLP). Crucially, the term $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ is fitted jointly. As for the linear model, the $\mathbf{\Sigma}\mathbf{\Sigma}^{\text{T}}$ is the noise covariance matrix and $\mathbf{W}$ is the standard Wiener process. In what follows we are going to often refer to this model as “neural model” or “neural emulator”.
@@
Note that this is the only Section in this work where we compare linear and nonlinear autoregressive models. The linear method is considered for two reasons. First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble-mean response to external perturbations. Second, and consistent with theoretical expectations [36], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings. This limitation restricts their usefulness for studying forced changes in distributions beyond the mean, which, dependent on the research question, can be a central objective in climate science.
@@ III.2.2 Stationary statistics
We integrate the two data-driven models inferred from data and specified in Eq. (6) and (7) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system. We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1. The neural model described in Equation (7) successfully captures the stationary distributions of all three modes $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$. In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of the first mode $x_{1}$. Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1. In summary, the neural emulator defined in Equation (7) is capable of capturing both the stationary statistics and temporal correlations of the full state vector $\mathbf{x}(t)$. At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@ III.2.3 Response of probability distributions to external perturbations
Impulse response functions
We now examine the impulse response of the original triad model (Equation (5)), the linear model (Equation (6)), and the nonlinear neural model (Equation (7)). Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$. The observables considered here are the mean and variance of the time-dependent perturbed distribution. The ensemble size considered here is $N_{e}=10^{5}$ trajectories. The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively. The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics. This is expected, as a linear model with a Gaussian invariant measure produces zero response to any quadratic functionals, such as the variance, consistent with the theoretical results in [36]. In contrast, the neural model captures the variance response with relatively high fidelity. As argued in Section II.2, the impulse response operator plays a foundational role in determining a model’s ability to capture the linear part of the response to general perturbations. We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response to forcing.
@@
Having evaluated the ability of the data-driven neural model to capture the impulse response operator, we now consider a more targeted test involving a specific, time-dependent forcing. We introduce a logarithmic forcing term to the right-hand side of the original triad model (Equation (5)), the linear model (Equation (6)), and the neural model (Equation (7)). The forcing is applied only to the first mode $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step. Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble-mean and ensemble-variance responses. Results are presented in Figure 3. The neural model accurately reproduces the forced responses across all three components. The linear model performs qualitatively well but exhibits significant deviations in the ensemble-mean response of the second mode $x_{2}$ as shown in Figure 3(b). This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)). This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings. As expected from the analysis of the response operator, the linear model is not able to capture the response in ensemble variance. The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in the second mode $x_{2}$. In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.

### v3

@@ II.1 Reduced-order stochastic climate models: theoretical and practical considerations
In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g. temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales. Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained spatiotemporal fields obtained via (i–iii). The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [67, 8], by retaining the first $n<N$ of such bases. The correspondent principal components (PCs), serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as in [68, 18, 69] may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [70, 71]. Identifying the “right” coarse-grained representation remains one of the hardest challenges in data-driven modeling, as no general mathematical principle exists to guide such choices in high-dimensional dynamical systems [43, 17, 17].
Inferring a stochastic model.
Given the coarse-grained data $\mathbf{x}$, the goal is to infer a stochastic Markovian model as in Eq. (2). We summarize several traditional approaches that motivate our work, while acknowledging that many other methods exist beyond those discussed here. The simplest example is the Linear Inverse Model (LIM) proposed by Penland [3], where the drift is linear and the noise is additive, both estimated directly from data. LIMs have shown forecasting skill for sea surface temperature anomalies comparable to climate models [4, 5, 72, 73], and extensions exist to handle colored noise [74]. Kravtsov et al. (2005) and Kondrashov et al. (2015) [6, 7] generalized this framework through multilevel regression, incorporating quadratic drift terms and non-Markovian effects. These models perform well for stationary statistics [8] and forecasting [75]. The work of Majda and Harlim extended the multilevel regression framework by proposing stability constraints [11]. A theoretically grounded alternative is the approach of Majda, Timofeyev, and Vanden-Eijnden [60, 61, 62, 63], dubbed MTV by the authors, which derives effective equations for the coarse-grained variables $\mathbf{x}$ rather than assuming them. MTV has been successful for simplified models, though less so for complex climate systems, likely due to choices of coarse-graining steps (i-iii) shown above rather than intrinsic limitations, see also discussion in [8]. We will expand on this approach in Section III and Appendix B.
Overall, these works demonstrated that robust reduced-order stochastic climate models can be built through relatively simple regression strategies. In this paper, we build on these insights and investigate how simple neural architectures can infer both the drift and state-dependent noise in Eq. (2), with minimal assumptions. Our neural models retain a decomposition into linear and nonlinear terms, $\mathbf{L}$ and $\mathbf{n}(\mathbf{x})$, as in the general representation of dynamical systems in Eq. (1). Furthermore, we will propose simple solutions to fit multiplicative (state-dependent) noise in high dimensions.
@@ II.2 Response theory: a (very) brief overview
In this paper, we argue that data-driven models should be tested on their ability to capture both stationary statistics and responses to external perturbations. Assessing responses to arbitrary forcings is ill-posed, as the space of possible perturbations is essentially infinite. A way forward is to consider small impulse perturbations, which allow us to define a response operator and derive general conclusions about causality and forced responses within the linear response regime.
@@
where $\langle\cdot\rangle$ refers to ensemble average and the subscript “p” refers to the perturbed system. In this paper we are going to focus on two different specific observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution. The two observables allow us to capture time-dependent responses in the mean and variance of the perturbed distribution respectively. The impulse response operator $\mathbf{R}(t)\in\mathbb{R}^{n,n}$ is then defined as:
@@
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77]. The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$ [78, 79, 17, 80, 18]. We refer to Appendix A for details on the relation between the response operator and causality.
@@
    $\delta\langle A(x_{k}(t))\rangle=\sum_{j}\int_{0}^{t}R_{k,j}(\tau)f_{j}(t-\tau)d\tau.$ (5)
A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom. Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation [18]. Moreover, since the linear response to small perturbations is given by the convolution formula in Eq. (5), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings. For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the response to small, general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well. Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [49], this is discussed in Appendix A.
A fundamental limitation of purely data-driven modeling of multiscale climate systems is that we cannot observe all degrees of freedom. Even when governing equations are known, as in meteorology, climate fields are only available at finite temporal and spatial resolution. Dynamics at smaller and faster scales are therefore missing from the data and often remain unparameterized in emulators, unlike in full climate models. An even greater challenge arises in many areas where we lack governing equations altogether to guide emulator construction. In both cases, data-driven approaches can mimic the observed variability of the retained variables, reproducing probability distributions, autocorrelation functions, and providing reasonable forecasts. Yet testing responses to perturbations, where unobserved dynamics becomes relevant, often exposes structural deficiencies that are otherwise hidden. The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of the Section meaningful for climate applications. The model was introduced by Majda and collaborators [46] and later used in [47] in the context of climate dynamics. In what follows, we present the model and refer to Appendix B for an in-depth discussion on its relevance as a building block for complex turbulent flows. The triad model considered in this study takes the following form:
@@ III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
This idealized system mimics the functional structure of turbulent geophysical flows, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities. The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics and Gaussian statistics. The quadratic interactions, here between $x_{1}$ and $x_{2}$, typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [16]. The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting. The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations. The model generates unimodal distributions with relatively small, but significant, deviations from Gaussianity, which is a relevant feature in the study of low-frequency climate variability [81] and found in both general circulation models and observational data, see e.g. [82, 83, 84].
III.2 The case of unknown equations but a fully observed state vector
@@
We consider the relatively simple case of a low-dimensional dynamical system with only one characteristic time scale (i.e. $\epsilon=1$). The parameters chosen for this first test are: $L_{11}=-2$, $L_{12}=1$, $L_{13}=1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=1$. We now integrate the system in (6) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units). Given the system’s decorrelation time scale (shown later), the resulting time series can be effectively treated as infinitely long. Given this long trajectory of the full state vector $\mathbf{x}(t)=[x_{1},x_{2},x_{3}](t)$, we learn data-driven models directly from the time series. We construct both linear and nonlinear stochastic models. We specify the functional forms of these models and assess their ability to reproduce both the stationary and perturbed statistics of the original triad system.
@@ III.2.1 Functional form of the data-driven models considered
We consider two data-driven models. To keep the main text focused, we leave the technical details of the fitting procedures, along with details of the neural networks, in Appendix C.
@@
    $\mathbf{L}$ is a linear operator and $\mathbf{n}(\mathbf{x})$ is parametrized by a Multilayer Perceptron (MLP). Crucially, the term $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ is fitted jointly. In what follows we are going to often refer to this model as “neural model” or “neural emulator”.
@@
Note that this is the only Section in this work where we compare linear and nonlinear autoregressive models. The linear method is considered for two reasons. First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations. Second, and consistent with theoretical expectations [47], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings. This limitation restricts their usefulness for studying forced changes in distributions beyond the mean.
@@ III.2.2 Stationary statistics
We integrate the two data-driven models inferred from data and specified in Eq. (7) and (8) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system. We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1. The neural model described in Equation (8) successfully captures the stationary distributions of $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$. In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of $x_{1}$. Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1. In summary, the neural emulator defined in Equation (8) is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$. At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@ III.2.3 Response of probability distributions to external perturbations
Impulse response functions.
We now examine the impulse response of the original triad model (Equation (6)), the linear model (Equation (7)), and the nonlinear neural model (Equation (8)). Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$. We choose as observables the mean and variance of the time-dependent perturbed distribution and refer to Appendix A for details on how to compute the response operator. The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively. The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics [47]. In contrast, the neural model captures the variance response with relatively high fidelity. We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response operator.
@@
we now consider a more targeted test involving a specific, time-dependent forcing. We introduce a logarithmic forcing term to the right-hand side of the original and data-driven triad model. The forcing is applied only to $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step. Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble mean and variance responses. Results are presented in Figure 3. The neural model (Eq. (8)) accurately reproduces the forced responses across all three components. The linear model (Eq. (7)) performs qualitatively well but exhibits significant deviations in the ensemble mean response of the second mode $x_{2}$ as shown in Figure 3(b). This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)). This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings. As expected, the linear model is not able to capture the response in ensemble variance. The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in $x_{2}$. In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.